rE=22% rD=__ rA=__ βE=__ βD=.625 βA=__ rf=10% rm=18% D/V=.8

What is βE ? (sample answer: 1.25)
What is rD ?  (sample answer: 12.5%)
What is rA ?  (sample answer: 12.5%)
What is βA ? (sample answer: 1.25)
Answer #1

rE=22%

re=rf+be*(rm-rf)
=>be=(22%-10%)/(18%-10%)=1.5

rd=10%+0.625*(18%-10%)=15.00%

rA=0.8*15%+(1-0.8)*22%=16.40%

Ba=0.8*0.625+(1-0.8)*1.5=0.8